OpenStudy (anonymous):

A store sells spherical helium-filled balloons. Each balloon contains 590 cubic inches of helium. What is the radius of the balloon?

OpenStudy (ybarrap):

Model the balloon as a sphere with radius r and volume V: $$ V=\frac 4 3 \pi r^3\implies r=\left (\frac{3V}{4\pi}\right )^{1/3}=\left (\frac{3(590)}{4\pi}\right )^{1/3} $$ That's it!

OpenStudy (ybarrap):

Do you have any questions?

OpenStudy (anonymous):

Nope, thanks :)

OpenStudy (anonymous):

wait, so I simplify the last step right?

OpenStudy (ybarrap):

You can if want. If you don't care about exact, just use a calculator.